OpenStudy (anonymous):

Vivien bought a box of apples for $10.50. She sold each apple for 60cents and made a profit of $1.50. How many apples were in the box?

OpenStudy (anonymous):

n number of apples. (n*0,60) - 10,50 = 1,50 solve this for n

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Let the number of apples be x.. So, she sold all apples at 0.60x.. And she bought apples at 10.50.. So, profit made = 0.60x - 10.50 As profit is given : 1.5 So, 0.60x - 10.50 = 1.5 0.60x = 12 x = 12/0.60 = 1200/60 = 20.. So. there were 20 apples in the box...